# DriveRX: A Vision-Language Reasoning Model for Cross-Task Autonomous Driving

### v1

@@ 4.1 Experiment Settings
Base Model: We adopt Align-DS-V[41] as our base model, which extends DeepSeek-R1-Distill-LLaMA-8B[29] with a CLIP-ViT-L/14-336 visual encoder [42], trained using the Align-Anything framework [43], and exhibits strong multi-modal reasoning performance.
Evaluated Models: We evaluate a diverse set of models, including the commercial GPT-4o (2024-08-06) [27], the general open-source model LLaVA-1.5 7B [44], LLaVA-NeXT 7B [45], InternVL3 8B/78B [46], Qwen2.5-VL 7B/72B [47], the reasoning MM-Eureka 7B [33], R1-Onevision 7B [37], and the domain-specific autonomous driving models DriveLM [35] and Dolphins [36].
Training Dataset: We use the DriveLM [35] dataset, which is constructed from nuScenes [48] and OpenLane-V2 [49] through a structured annotation process. Following the AutoDriveRL data selection strategy, we compute frame-level scores $D_{i}$ using Align-DS-V and retain samples within the range $[25,45]$. To ensure a non-overlapping evaluation, all samples appearing in DriveBench [35] are explicitly excluded. From the remaining subset, we randomly sample 6K instances for training. See Appendix C for the dataset composition.
Reward: We first apply a rule-based filter that assigns a reward of 0 to responses exceeding 4K tokens or exhibiting substantial repetition. The remaining responses are then evaluated using task-specific prompts by Qwen2.5-72B-Instruct [28] (see Appendix I.1). The consistency between human judgments and model scores is analyzed in Appendix F.
Evaluation: We evaluate on DriveBench [35], which spans four driving tasks and consists of a clean set with re-sampled questions and a corruption set featuring 15 types of visual degradation (see Appendix J) for robustness testing. We additionally evaluate on DriveLM-Hard, a challenging subset selected using the AutoDriveRL data selection strategy, where frame-level scores $D_{i}$ (defined in Eq. 1) fall within the range $[10,31]$. The $D_{i}$ scores are computed by averaging model scores from Align-DS-V, Qwen2-VL-7B, and Qwen2-VL-72B.
Metrics: We follow DriveBench evaluation prompts and use GPT-3.5-Turbo as the judge model, with GPT score as the evaluation metric. Following prior work [35, 50, 51, 52, 53], we discard BLEU [54] and ROUGE [55], as they were designed for translation and are not suitable for evaluating decision quality in autonomous driving.
For more experimental setup details, please refer to Appendix K.
@@ 4.2 Main Result
DriveRX shows strong overall performance and achieves the best result in behavior tasks: As shown in Figure 1, DriveRX demonstrates competitive performance across all four tasks. On the clean test sets of Perception, Prediction, and Planning, it performs closely to GPT-4o and the larger 70B-scale models, despite its smaller size. Notably, it achieves the highest score ($\mathbf{62.02}$) on the critical Behavior task, surpassing all baselines including the closed-source GPT-4o ($55.04$), the open-source Qwen2.5-VL 72B ($55.57$), the reasoning model MM-Eureka ($50.50$), and the domain-specific DriveLM ($42.78$). These results highlight the effectiveness of the AutoDriveRL framework, as improvements in earlier tasks lead to significant gains in the final behavior task.
@@
GPT-4o [27] - Commercial $41.51$ $45.39$ $52.28$ $50.92$ $84.82$ $84.43$ $55.04$ $53.97$
LLaVA-1.5 [44] $7$ B Open $23.22$ $22.95$ $22.02$ $17.54$ $29.15$ $31.51$ $13.60$ $13.62$
LLaVA-NeXT [45] $7$ B Open $24.15$ $19.62$ $35.07$ $35.89$ $45.27$ $44.36$ $48.16$ $39.44$
Qwen2.5-VL [28] $7$ B Open $31.92$ $27.16$ $42.26$ $47.13$ $55.70$ $57.97$ $47.40$ $42.82$
InternVL3 [46] $8$ B Open $33.21$ $\mathbf{33.87}$ $41.21$ $38.93$ $70.97$ $56.71$ $51.47$ $50.32$
Qwen2.5-VL [28] $72$ B Open $\mathbf{37.56}$ $32.25$ 54.89 50.47 $77.18$ $74.01$ $\mathbf{55.57}$ $51.65$
InternVL3 [46] $78$ B Open 39.44 35.62 $50.93$ $\mathbf{49.71}$ 82.24 78.41 $50.73$ $\mathbf{52.03}$
MM-Eureka [33] $7$ B Reasoning $34.44$ $28.92$ $40.31$ $41.72$ $65.38$ $60.18$ $49.4$ $47.59$
R1-Onevision [37] $7$ B Reasoning $28.64$ $25.40$ $51.22$ $46.91$ $55.12$ $55.54$ $41.57$ $39.88$
DriveLM [35] $7$ B Specialist $16.85$ $16.00$ $44.33$ $39.71$ $68.71$ $67.60$ $42.78$ $40.37$
Dolphins [36] $7$ B Specialist $9.59$ $10.84$ $32.66$ $29.88$ $52.91$ $53.77$ $8.81$ $8.25$
Align-DS-V [41] $8$ B Reasoning $31.41$ $27.78$ $41.51$ $37.86$ $51.16$ $57.66$ $50.53$ $40.98$
DriveRX $8$ B Reasoning $32.88$ $29.92$ $\mathbf{52.20}$ $48.29$ $\mathbf{78.63}$ $\mathbf{75.53}$ 62.02 55.24
DriveRX maintains robust performance under visual corruptions: In Figure 1, DriveRX exhibits consistently strong performance across all tasks under the corruption setting, demonstrating resilience to various types of visual degradation. Notably, it achieves the highest score of $\mathbf{55.24}$ on the Behavior task, outperforming all other models. This result highlights the robustness of DriveRX in handling distributional shifts, and further demonstrates the benefit of RL: the model receives step-wise feedback during training, enabling reliable behavior decision-making even under degraded visual conditions.
DriveRX achieves stable performance in complex traffic conditions: As shown in Figure 2, DriveRX performs strongly on DriveLM-Hard, a challenging test split we construct by selecting high-difficulty frames from the original DriveLM dataset. This benchmark features more congested traffic, diverse agents, and denser interactions across the same four tasks (see Appendix H). These results demonstrate that DriveRX retains reliable performance in complex and high-uncertainty scenarios, confirming its robustness and generalization. Such capabilities are critical for real-time autonomous driving applications that demand consistent reasoning under diverse and dynamic conditions.
@@
GPT-4o [27] - Commercial $36.12$ $65.90$ $51.42$ $42.00$
Qwen2.5-VL [47] $7$B Open $26.10$ $60.13$ $32.41$ $19.86$
InternVL3 [46] $8$B Open $31.29$ $38.95$ $31.78$ $30.48$
Qwen2.5-VL [47] $72$B Open $32.49$ $58.63$ $\mathbf{50.55}$ $40.02$
InternVL3 [46] $78$B Open $32.98$ $58.67$ $45.82$ $35.69$
MM-Eureka [33] $7$B Reasoning $29.09$ $\mathbf{67.41}$ $29.67$ $23.21$
R1-Onevision [37] $7$B Reasoning $25.51$ $46.60$ $34.45$ $29.03$
DriveLM [35] $7$B Specialist $28.78$ $19.93$ $22.60$ $30.75$
Align-DS-V [41] $8$B Reasoning $29.45$ $52.90$ $29.15$ $32.48$
DriveRX $8$B Reasoning $34.83$ $71.84$ $51.42$ $\mathbf{36.82}$
5 Analysis
5.1 Training Analysis
We analyze the training dynamics of DriveRX under AutoDriveRL to understand cross-task optimization. In Figure 2(a), the overall training remains stable, with steadily increasing rewards. This suggests that our task-specific reward models are effective in guiding learning across heterogeneous tasks. In Figure 2(b), the reward curves exhibit a clear staged trend: Perception and Prediction improve rapidly in early training and quickly saturate, while Planning and Behavior progress more slowly and gradually. This pattern reflects the structural hierarchy of the tasks—lower-level tasks are relatively simple, focusing on visual understanding and pattern recognition, whereas higher-level tasks require multi-step reasoning and rely on semantic representations generated by preceding tasks.
We also track the average response length during training (Figure 2(c)). In the early stages, response length increases, indicating that the model explores more expressive outputs. As training progresses and reward signals stabilize, responses gradually become shorter and more concise. Figure 2(d) shows the ratio of clipped responses exceeding the maximum length threshold, which closely follows the trend of response length—higher lengths lead to more clipping, while this ratio decreases as the model converges. These results confirm the effectiveness of our length-penalizing reward design, which encourages the model to reduce redundancy while maintaining output quality.
@@ 5.1 Training Analysis
Figure 2: Training dynamics of DriveRX over reinforcement learning steps.
5.2 Impact of Vision Encoders on Perception Tasks
Second, the visual encoder plays a critical role in determining the upper bound of perception task performance during training. In our experiments, the improvement of DriveRX over its base model Align-DS-V is limited ($+1.47$), despite a substantial gain on the behavior task ($+11.49$). As shown in the training curves (Figure 2(b)), the perception reward for DriveRX saturates early (within 50 steps) and remains below 50, indicating a clear performance ceiling. We attribute this to the outdated architecture of CLIP-ViT-L/14, an early-generation ViT model whose limited perceptual capability falls behind modern ViT-based encoders.
However, high-performing models like Qwen2.5-VL [28] and InternVL3 [46] typically adopt dynamic resolution mechanisms, which allow them to process high-resolution inputs more effectively. While this improves visual understanding and perception scores, it also leads to significantly longer image token sequences. For instance, Qwen2.5-VL generates substantially more vision tokens per image compared to fixed-resolution VLMs like Align-DS-V. This increase in token length introduces notable inference latency, posing challenges for real-time decision-making in autonomous driving. In contrast to these models, DriveRX incorporates a fixed-length feature processing pipeline in its architectural design. While this design sacrifices some perceptual performance due to the output dimensionality constraints of the CLIP encoder, it achieves three critical advantages for practical deployment: deterministic inference efficiency, faster processing speed, and more predictable inference behavior.
Model Size [Uncaptioned image] Vision Encoder [Uncaptioned image] Response Length [Uncaptioned image] Image Tokens [Uncaptioned image] Perception ($\uparrow$) [Uncaptioned image] Infer times
[…]
Table 3: Impact of different vision encoders on the perception task. We report DriveBench perception scores (higher is better), average response length (in tokens) per question, number of image tokens per image, and inference time per question (in seconds). For a fair comparison, inference time is reported only among models of the same size.
@@ 5.3 Case Study
Perception: Align-DS-V focuses on describing the overall scene—e.g., “a rainy street scene with a vehicle on the left side, partially obscured by fog or mist”—which aligns more with a captioning-style response. In contrast, DriveRX pays closer attention to driving-relevant visual cues, such as the wet road surface and the current pose of the vehicle.
@@
Prediction: Align-DS-V erroneously concludes that the vehicle is “aligned with the lane, indicating it may be driving straight”, overlooking the subtle orientation bias toward the right. This leads to a misprediction of the vehicle’s intent. DriveRX correctly identifies the vehicle as angled across the lane, indicating a right turn, and additionally predicts its motion status (e.g., speed and steering).
Planning: Align-DS-V adopts an elimination-based reasoning strategy—e.g., “lack of visible turning signals, the vehicle is not steering.” This VQA-style approach relies solely on explicit visual cues and therefore struggles to capture the latent intent embedded in dynamic context. In contrast, DriveRX integrates implicit cues from the surrounding context with prior outputs from the perception and prediction stages, enabling more robust planning decisions—particularly in cases where signals are incomplete or the environment is visually degraded.
@@
Behavior: As a result of the flawed intermediate reasoning, Align-DS-V makes an incorrect decision (“going straight”), whereas DriveRX produces the correct and interpretable decision to steer right, grounded in structured multi-stage analysis.

### v2

@@ 4.1 Experiment Settings
Base Model: We adopt Align-DS-V(PKU-Alignment, 2025) as our base model, which extends DeepSeek-R1-Distill-LLaMA-8B(Guo et al., 2025) with a CLIP-ViT-L/14-336 visual encoder (Radford et al., 2021), trained using the Align-Anything framework (Ji et al., 2024), and exhibits strong multi-modal reasoning performance.
Evaluated Models: We evaluate a diverse set of models, including the commercial GPT-4o (2024-08-06) (OpenAI, 2024a), the general open-source model LLaVA-1.5 7B (Liu et al., 2024a), LLaVA-NeXT 7B (Liu et al., 2024b), InternVL3 8B/78B (Zhu et al., 2025), Qwen2.5-VL 7B/72B (Bai et al., 2025), the reasoning MM-Eureka 7B (Meng et al., 2025), R1-Onevision 7B (Yang et al., 2025b), and the domain-specific driving models DriveLM (Sima et al., 2024) and Dolphins (Ma et al., 2024).
Training Dataset: We use the DriveLM (Sima et al., 2024) dataset, which is constructed from nuScenes (Caesar et al., 2020) and OpenLane-V2 (Wang et al., 2023) through a structured annotation process. Following the AutoDriveRL data selection strategy, we compute frame-level scores $D_{i}$ using Align-DS-V and retain samples within the range $[25,45]$. To ensure a non-overlapping evaluation, all samples appearing in DriveBench (Sima et al., 2024) are explicitly excluded. From the remaining subset, we randomly sample 6K instances for training. See Appendix C for the dataset composition.
Reward: We first apply a rule‑based filter that assigns a reward of 0 to responses exceeding 4K tokens or exhibiting substantial repetition. The remaining responses are then evaluated using task‑specific prompts by Qwen2.5‑72B‑Instruct (Team, 2024) (see Appendix M.1). The consistency between human judgments and model scores, as well as the scoring latency, is analyzed in Appendix I and Appendix O.3, respectively.
Evaluation: We evaluate on DriveBench (Sima et al., 2024), which spans four driving tasks and consists of a clean set with re-sampled questions and a corruption set featuring 15 types of visual degradation (see Appendix N) for robustness testing. We additionally evaluate on DriveLM-Hard, a challenging subset selected using the AutoDriveRL data selection strategy, where frame-level scores $D_{i}$ (defined in Eq. 1) fall within the range $[10,31]$. The $D_{i}$ scores are computed by averaging model scores from Align-DS-V, Qwen2-VL-7B, and Qwen2-VL-72B.
Metrics: We follow DriveBench evaluation prompts and use GPT-3.5-Turbo as the judge model, with GPT score as the evaluation metric. Following prior work (Sima et al., 2024; Tran et al., 2019; Anderson et al., 2016; Akter et al., 2022; Evtikhiev et al., 2023), we discard BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004), as they were designed for translation and are not suitable for evaluation in autonomous driving. For more experimental setup and metrics details in Appendix O.
@@ 4.2 Main Result
DriveRX shows strong overall performance and achieves the best result in behavior tasks: As shown in Table 1, DriveRX demonstrates competitive performance across all four tasks. On the clean test sets of Perception, Prediction, and Planning, it performs closely to GPT-4o and the larger 70B-scale models, despite its smaller size. Notably, it achieves the highest score ($\mathbf{62.02}$) on the critical Behavior task, surpassing all baselines including the closed-source GPT-4o ($55.04$), the open-source Qwen2.5-VL 72B ($55.57$), the reasoning model MM-Eureka ($50.50$), and the domain-specific DriveLM ($42.78$). These results highlight the effectiveness of the AutoDriveRL framework, as improvements in earlier tasks lead to significant gains in the final behavior task.
@@
GPT-4o (OpenAI, 2024a) - Commercial $41.51$ $45.39$ $52.28$ $50.92$ $84.82$ $84.43$ $55.04$ $53.97$
LLaVA-1.5 (Liu et al., 2024a) 77 B Open $23.22$ $22.95$ $22.02$ $17.54$ $29.15$ $31.51$ $13.60$ $13.62$
LLaVA-NeXT (Liu et al., 2024b) 77 B Open $24.15$ $19.62$ $35.07$ $35.89$ $45.27$ $44.36$ $48.16$ $39.44$
Qwen2.5-VL (Team, 2024) 77 B Open $31.92$ $27.16$ $42.26$ $47.13$ $55.70$ $57.97$ $47.40$ $42.82$
InternVL3 (Zhu et al., 2025) 88 B Open $33.21$ $\mathbf{33.87}$ $41.21$ $38.93$ $70.97$ $56.71$ $51.47$ $50.32$
Qwen2.5-VL (Team, 2024) $72$ B Open $\mathbf{37.56}$ $32.25$ 54.89 50.47 $77.18$ $74.01$ $\mathbf{55.57}$ $51.65$
InternVL3 (Zhu et al., 2025) $78$ B Open 39.44 35.62 $50.93$ $\mathbf{49.71}$ 82.24 78.41 $50.73$ $\mathbf{52.03}$
MM-Eureka (Meng et al., 2025) 77 B Reasoning $34.44$ $28.92$ $40.31$ $41.72$ $65.38$ $60.18$ $49.40$ $47.59$
R1-Onevision (Yang et al., 2025b) 77 B Reasoning $28.64$ $25.40$ $51.22$ $46.91$ $55.12$ $55.54$ $41.57$ $39.88$
DriveLM (Sima et al., 2024) 77 B Specialist $16.85$ $16.00$ $44.33$ $39.71$ $68.71$ $67.60$ $42.78$ $40.37$
Dolphins (Ma et al., 2024) 77 B Specialist $9.59$ $10.84$ $32.66$ $29.88$ $52.91$ $53.77$ $8.81$ $8.25$
Align-DS-V (PKU-Alignment, 2025) 88 B Reasoning $31.41$ $27.78$ $41.51$ $37.86$ $51.16$ $57.66$ $50.53$ $40.98$
DriveRX 88 B Reasoning $32.88$ $29.92$ $\mathbf{52.20}$ $48.29$ $\mathbf{78.63}$ $\mathbf{75.53}$ 62.02 55.24
DriveRX maintains robust performance under visual corruptions: In Table 1, DriveRX exhibits consistently strong performance across all tasks under the corruption setting, demonstrating resilience to various types of visual degradation. Notably, it achieves the highest score of $\mathbf{55.24}$ on the Behavior task, outperforming all other models. This result highlights the robustness of DriveRX in handling distributional shifts, and further demonstrates the benefit of RL: the model receives step-wise feedback during training, enabling reliable decision-making even under visual corruptions.
DriveRX achieves stable performance in complex traffic conditions: As shown in Table 2, DriveRX performs strongly on DriveLM-Hard, a challenging test split we construct by selecting high-difficulty frames from the DriveLM dataset. This benchmark features congested traffic, diverse agents, and denser interactions across the same four tasks (see Appendix L). These results demonstrate that DriveRX maintains reliable performance in complex and high-uncertainty scenarios, confirming its robustness and generalization. Such capabilities are critical for safety-critical autonomous driving applications that demand consistent reasoning under diverse and dynamic conditions.
@@
GPT-4o (OpenAI, 2024a) - Commercial $36.12$ $65.90$ $51.42$ $42.00$
Qwen2.5-VL (Bai et al., 2025) 77B Open $26.10$ $60.13$ $32.41$ $19.86$
InternVL3 (Zhu et al., 2025) 88B Open $31.29$ $38.95$ $31.78$ $30.48$
Qwen2.5-VL (Bai et al., 2025) 7272B Open $32.49$ $58.63$ $\mathbf{50.55}$ $40.02$
InternVL3 (Zhu et al., 2025) 7878B Open $\mathbf{32.98}$ $58.67$ $45.82$ $35.69$
MM-Eureka (Meng et al., 2025) 77B Reasoning $29.09$ $\mathbf{67.41}$ $29.67$ $23.21$
R1-Onevision (Yang et al., 2025b) 77B Reasoning $25.51$ $46.60$ $34.45$ $29.03$
DriveLM (Sima et al., 2024) 77B Specialist $28.78$ $19.93$ $22.60$ $30.75$
Align-DS-V (PKU-Alignment, 2025) 88B Reasoning $29.45$ $52.90$ $29.15$ $32.48$
DriveRX 88B Reasoning $34.83$ $71.84$ $51.42$ $\mathbf{36.82}$
Overall, our results on DriveBench and DriveLM-Hard demonstrate that DriveRX consistently produces stable and interpretable stage-wise reasoning across challenging, corrupted, and long-tail scenarios. These findings confirm that DriveRX is well suited as a high-level semantic reasoning backbone for autonomous driving: its structured Perception→Prediction→Planning→Behavior chain provides reliable cross-task semantics and intention-level understanding.
5 Training Analysis
We analyze the training dynamics of DriveRX under AutoDriveRL to understand cross-task optimization. In Figure 2(d)(a), the overall training remains stable, with steadily increasing rewards. This suggests that our task-specific reward models are effective in guiding learning across heterogeneous tasks. In Figure 2(d)(b), the reward curves exhibit a clear staged trend: Perception and Prediction improve rapidly in early training and quickly saturate, while Planning and Behavior progress more slowly and gradually. This pattern reflects the structural hierarchy of the tasks—lower-level tasks are relatively simple, focusing on visual understanding and pattern recognition, whereas higher-level tasks require multi-step reasoning and rely on semantic representations generated by preceding tasks.
We also track the average response length during training (Figure 2(d)(c)). In the early stages, response length increases, indicating that the model explores more expressive outputs. As training progresses and reward signals stabilize, responses gradually become shorter and more concise. Figure 2(d)(d) shows the ratio of clipped responses exceeding the maximum length threshold, which closely follows the trend of response length—higher lengths lead to more clipping, while this ratio decreases as the model converges. These results confirm the effectiveness of our length-penalizing reward design, which encourages the model to reduce redundancy while maintaining output quality.
@@ 5 Training Analysis
Figure 2: Training curves of DriveRX over reinforcement learning steps.
@@ 5.1 Case Study
Perception: Align-DS-V mainly provides scene-level descriptions (e.g., “a rainy street scene with a vehicle on the left”), while DriveRX focuses on driving-critical cues like the wet road surface and vehicle pose.
@@
Prediction: Align-DS-V overlooks subtle orientation bias and mispredicts the vehicle’s intent as going straight. In contrast, DriveRX correctly infers a right turn and estimates motion status (e.g., speed, steering).
Planning: Relying solely on explicit visual evidence (e.g., “no turn signal”), Align-DS-V struggles under incomplete cues. DriveRX integrates implicit context with earlier outputs, enabling more reliable planning in degraded environments.
@@
Behavior: Due to its flawed reasoning chain, Align-DS-V makes an incorrect “go straight” decision, whereas DriveRX outputs the correct and interpretable action to steer right.
6 Applications
We explore two applications where DriveRX serves as a high-level reasoning backbone that provides semantic supervision for downstream planning/control models: trajectory prediction and action generation. These applications illustrate how a reasoning-enhanced VLM can support cross-task semantic understanding and intention reasoning, and how its structured outputs can effectively guide downstream models in complex driving scenarios. Additional details are provided in Appendix H.
Table 3: Open-loop testing on nuScenes.
[…]
Command Mean 4.57 5.71
BLIP-RT-2 2.63 2.77
DriveLM-Agent 1.39 1.67
DriveRX-Agent 1.53 1.12
DriveRX-Agent. Following prior work (Zitkovich et al., 2023; Sima et al., 2024), we extend the vocabulary of DriveRX by discretizing the trajectory coordinates (the waypoints of the ego vehicle) into 512 bins, empirically determined based on the statistics of the training set trajectories. We then redefine the tokens in the DriveRX language tokenizer, creating distinct tokens for each bin, and fine-tune the VLM using this redefined vocabulary. This enables DriveRX-Agent to directly generate the corresponding vehicle trajectory. We conduct open-loop evaluation on the nuScenes dataset, computing the ADE and Collision Rate (Col.) by averaging the values at the 1s, 2s, and 3s time horizons, following the evaluation settings in DriveLM (Sima et al., 2024), and the results in Table 3 show that DriveRX-Agent outperforms the baseline DriveLM-Agent, highlighting how improved reasoning in VLMs can directly translate into better trajectory prediction.
DriveRX-VLA. We further distill reasoning data from DriveRX and use supervised fine-tuning to train a smaller Vision-Language-Action(VLA) model based on the Qwen2.5-VL-3B (Bai et al., 2025). The prompt set is derived from the DriveLM-CARLA (Sima et al., 2024). Continuous vehicle trajectories are discretized into a series of physical action tokens using a K-means clustering approach, forming an action codebook with 2048 discrete tokens representing short-term spatial position changes and heading movements (steer, throttle, brake). These tokens are integrated into the VLM vocabulary, allowing the model to directly predict control actions during inference. Trained on reasoning data from DriveRX, the resulting model is capable of generating appropriate actions even in complex driving scenarios. Closed-loop evaluations on the Bench2Drive (Jia et al., 2024) benchmark (see Table 4, the planning frequency is set to 2 Hz) show that DriveRX-VLA achieves competitive performance. In the evaluation, While DriveRX-VLA slightly lags behind the top-performing model, it nonetheless achieves competitive results using significantly less training data compared to Orion, thereby demonstrating the efficacy of reasoning-based distillation.
Overall, DriveRX provides stable and interpretable reasoning across four tasks, enabling it to serve as an effective high-level semantic backbone for downstream driving tasks. Its structured outputs offer high-quality supervisory signals—both as reasoning traces and reward feedback—that enhance trajectory prediction and action generation when distilled into smaller models. The results of DriveRX-Agent and DriveRX-VLA confirm the feasibility of leveraging a unified reasoning-enhanced VLM to improve downstream planning and control models.
Table 4: Closed-loop Results on the Bench2Drive.
Methods Driving Score $\uparrow$ Success Rate (%) $\uparrow$ Efficiency $\uparrow$ Comfortness $\uparrow$
TCP-traj (Wu et al., 2022) 59.90 30.00 76.54 18.08
AD-MLP (Zhai et al., 2023) 18.05 0.00 48.45 22.63
UniAD-Base (Hu et al., 2023) 45.81 16.36 129.21 43.58
VAD (Jiang et al., 2023) 42.35 15.00 157.94 46.01
DriveTransformer-Large (Jia et al., 2025) 64.22 33.08 70.22 16.01
Orion (Fu et al., 2025) 77.74 54.62 151.48 17.38
DriveRX-VLA 72.36 45.23 138.47 17.36